A plant grew 10 centimeters in 4 months. It is now 0.8 m tall. How tall was
the plant 3 months ago?

Answer by ikleyn(52878)

.
The rate of growth is 10/4 = 2.5 cm per month.


The height of the plant 3 months ago was  80 cm - 3*2.5 cm = (80 - 7.5) cm = 72.5 cm.

Answered, explained and completed.